How can we know oxidation state of P in Ba(H2PO2)2 please explain

Grade:11

1 Answers

Arun
6 years ago
Dear Chandan
 
the ions with one or two oxidation number(basic)u hav 2 memorise or know it!!!! 
example: 
Ba,Barium=+2 
O,oxygen=-2 
H,hydrogen=+1 

the oxidation number of whole molecules=0 
put the phophurus as X 

+2+2(2(+1)+X+(2)-2)=0 
2+4+2X-8=0 
2X=8-4-2 
=2 
X=+1 

the oxidation number is +1
 
 
 
Regards
Arun (askIITians forum expert)